# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

### v2

@@ 17. Clifford Algebras: Clifford and Lie algebra
Let’s look at the first few of these. For ${\mathcal{C}}_{0}$ we are handed $\mathbb{R}$} itself. The algebra ${\mathcal{C}}_{1}$ is given by additively by two copies of $\mathbb{R}$}, one for $1$ and one for $e_{1}$, with $e_{1}^{2}=-1$ and is easily seen to be the complex numbers $\mathbb{C}$}. For ${\mathcal{C}}_{2}$ we have additive generators $1$, $e_{1}$, $e_{2}$ and $e_{1}e_{2}$, and
$(e_{1}e_{2})^{2}=e_{1}e_{2}e_{1}e_{2}=-(e_{1}e_{1})(e_{2}e_{2})=-1$
and this is just the quaternions $\mathbb{H}$}.
@@
    ${\mathcal{C}}_{k+8}\,\cong\,{\mathcal{C}}_{k}\otimes M_{16}({\mathbb{R}})\qquad\qquad{\mathcal{C}}^{c}_{k+2}\,\cong\,{\mathcal{C}}^{c}_{k}\otimes M_{2}({\mathbb{C}})$
@@
Each Clifford algebra $A$ has a Lie group $Lie(A)\subset A$ associated to it, namely the elements in it that are “unitary” in the sense that $aa^{*}=a^{*}a=1$. We write
@@
We may realize the the Lie groups ${\mathcal{L}}_{j}$ of the Clifford algebras in a different way. Fix some large index $t$ and integer $k$. Fix a faithful $C^{*}$-representation
@@
$\big<O_{16t},\,\,\,J_{1},\dots J_{n}\big$>
denote the subgroup of $O_{16t}$ consisting of all of the matrices that commute with $J_{1},\dots J_{n}$. Then we have the following result.
@@
$U_{8t}=\big<O_{16t},\,\,\,J_{1}\big$>
$Sp_{4t}=\big<O_{16t},\,\,\,J_{1},J_{2}\big$>
$Sp_{2t}\times Sp_{2t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3}\big$>
$Sp_{2t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4}\big$>
$U_{2t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5}\big$>
$O_{2t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6}\big$>
$O_{t}\times O_{r}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6},J_{7}\big$>
$O_{t}=\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6},J_{7},J_{8}\big$>
The sequence may be extended infinitely in each direction when $t$ is a sufficiently large power of $2$. The analogous result holds in the complex setting as well.
@@
$O_{1}\hookrightarrow U_{1}\hookrightarrow Sp_{1}\hookrightarrow Sp_{1}\times Sp_{1}\hookrightarrow Sp_{2}\hookrightarrow U_{4}\hookrightarrow O_{8}\hookrightarrow O_{8}\times O_{8}\hookrightarrow O_{16}\qquad(**)$
@@
$O_{16}\hookleftarrow U_{8}\hookleftarrow Sp_{4}\hookleftarrow Sp_{2}\times Sp_{2}\hookleftarrow Sp_{2}\hookleftarrow U_{2}\hookleftarrow O_{2}\hookleftarrow O_{1}\times O_{1}\hookleftarrow O_{1}\qquad(@@)$
@@
from (@@). So the groups appear in opposite order. What is really neat is that the maps are the same - just re-ordered! We have inserted the numbering of the maps simply to facilitate comparison. The same thing happens in the complex setting, but there are only two groups involved, and so it passes without notice. We want to refer to these two different points of view.
@@ 18. Clifford Algebras and Bott Periodicity: the Real Case
    There is a weak homotopy equivalence
@@
    Taking $A={\mathbb{R}}$ and $r=7$ in the previous result, there is a weak homotopy equivalence
@@
Please note that this strongly ties Clifford algebras, or, more precisely, their Lie groups, to the Homotopy Road Map. Rather than regarding the first column as coming to us from Bott Periodicity, we may think of the column as simply presenting yet a different look at the Clifford and Lie algebra perspectives.
@@ 19. Clifford Algebras and Bott Periodicity: the Complex Case
Suppose that $A$ is a unital complex Banach algebra.
Here’s some notation that we will use in this section:
@@
    Taking $A={\mathbb{C}}$ there is a weak homotopy equivalence
@@ 20. The Basics: Symmetric Spaces and a Timely Example
Suppose that $M$ is a connected Riemann manifold and let $p\in M$. There is an exponential map $exp_{p}:T_{p}M\to M.$ We may choose a small open ball $B_{p}$ about the origin of $T_{p}M$ and label its image as $V_{p}\equiv exp_{p}(B_{p})\subset M.$ If $B_{p}$ is small enough then the map
$exp_{p}:B_{p}\longrightarrow V_{p}$
@@
$V_{p}\overset{exp_{p}^{-1}}{\longrightarrow}B_{p}\overset{u\to-u}{\longrightarrow}B_{p}\overset{exp_{p}}{\longrightarrow}V_{p}.$
@@
A Lie algebra $\mathfrak{g}$ over a field $\mathbb{F}$} is an $\mathbb{F}$}-vector space together with a map
@@
For example, take $\mathfrak{g}=M_{k}({\mathbb{F}})$ with [A,B]=ABBA[A,B]=AB-BA.
@@
Now, what is the relationship between Lie groups and Lie algebras. Here are the facts:
@@
    If $G$ is a Lie group then $Lie(G)$ is a Lie algebra, with $dim(G)$ (as a compact smooth manifold) the same as $dim(Lie(G))$ (as a real vector space.)
@@
    If we restrict attention to simply connected797979i.e. with $\pi_{1}(G)=0$. Lie groups, then there is a 1-1 correspondence
@@ 20.1. Topology and Physics, by Nadav Orion
for the joint system (both spaces with the usual inner product). A typical Hamiltonian for such a system contains both terms that affect a single qubit ($\sigma_{j}\otimes I$ and $I\otimes\sigma_{k}$) and terms that describe interactions between the two ($\sigma_{j}\otimes\sigma_{k}$), such that the set of Hamiltonians is all of the Hermitian $4\times 4$ matrices, which is the Lie algebra $\mathfrak{u}(4)$. The evolution operator space is thus $U_{4}$, a well-known compact complex symmetric space.
@@ 21. Quantum Symmetry
Suppose that $G$ is a compact Lie group and $K$ is a closed sub-Lie group in $G$. Then the Lie algebra of $G$ decomposes as real vector spaces
@@
There are, however, other possibilities, the first of which being time-reversal symmetry $T$. Possessing this symmetry means that the system would look the same, even at the atomic level, if time were reversed. So in general we can write time-reversal symmetry as $T$ with $T^{2}=I$ or $T^{2}=-I$ and the custom is to write 0 if it is not present. The Timely Example was indeed an example of time-reversal symmetry with $T^{2}=I$.
@@
Finally, it is possible to have a system that has neither time reversal nor charge conjugation symmetry, but it’s symmetrical under the combination of time reversal and charge conjugation. We also may have $S=CT$ with $S^{2}=1$ .
@@ C BdG 0 -1 0 $\mathbf{Sp}\,\,\,\cong\,\,\,\Omega^{4}\mathbf{O}$ ${\mathcal{R}}_{5}$ 0
we see at once that $KO$-theory and $K$-theory are intimately related to this story. The homotopy properties of the ${\mathcal{R}}_{n}$ are detected by $KO$ and hence connected to real vector bundles. This allows topological properties of the various symmetries to be detected and utilized. Much current physics research in topological insulators and related matters uses $KO^{*}$, and this was one of the main motivations for writing this work .
@@
22. Real and Complex $K$-theory
@@
    The product structure on the ring $KO_{*}({\mathbb{R}})$.
@@
    The map $c\colon KO_{*}(A)\rightarrow K_{*}(A_{\mathbb{C}})$ is defined as in Section 6.
@@
    But this projection is unitarily equivalent to $\left(\begin{smallmatrix}a&-b\\ b&a\end{smallmatrix}\right)$, by conjugating by $u=\left(\begin{smallmatrix}-1&0\\ 0&-1\end{smallmatrix}\right)$. It follows that $r\psi[p]=r[p]$.
@@
That is, $KO_{*}({\mathbb{R}})$ is the quotient of the graded polynomial ring over the integers generated by $\eta,\xi,\beta_{0},$ and $\beta_{0}^{-1}$ modulo the ideal generated by the relations specified above.
@@
$\cdots\rightarrow KO_{n}(A)\xrightarrow{~\eta~}KO_{n+1}(A)\xrightarrow{~c~}K_{n+1}(A_{\mathbb{C}})\xrightarrow{~r\beta^{-1}}KO_{n-1}(A)\rightarrow\cdots$ (3)
@@
Consider the map $f_{*}$ applied to the short exact sequence of Equation (3).
@@
If $KO_{n}(A)=0$ for all $n$, then the long exact sequence (3) immediately implies that $K_{n}(A_{\mathbb{C}})=0$ for all $n$. Conversely, suppose that $K_{n}(A_{\mathbb{C}})=0$ for all $n$, then the long exact sequence (3) implies that $\eta_{n}$ is an isomorphism for all $n$. However, the relation $\eta^{3}=0$ then implies that $KO_{n}(A)=0$ for all $n$. ∎
@@
    If $A$ is a real $C^{*}$-algebra with complexification $A\overset{c}{\longrightarrow}A_{\mathbb{C}}$, then the map $c$ induces a natural transformation of spectral sequences
    $E_{p,q}^{r}(A)\overset{c}{\longrightarrow}E_{p,q}^{r}(A_{\mathbb{C}}).$
@@
Suppose that $X$ is a finite simplicial complex838383imagine a CW-complex with finitely many cells attached with skeleta [^84] For CW complexes, let $X_{0}$ be the finite set of points you start with, $X_{1}$ is what you have after adding all of the one-dimensional cells, etc. $X_{j}$. Let $\{x_{o}\}$ denote the space with a single point. Then there is a spectral sequence $E_{r}^{p,q}$ ($r\geq 1$, $\infty<p,q<+\infty$), with
$E_{1}^{p,q}\cong C^{p}(X;K^{q}(\{x_{o}\})$
($p$- cochains taking values in $K^{q}(\{x_{o}\}$ ), with $d^{1}$ usual coboundary operator, and
$E_{2}^{p,q}\cong H^{p}(X;K^{q}(\{x_{o}\})$
$E_{\infty}^{p,q}\cong G_{p}K^{p+q}(X).$
where $G_{p}K^{p+q}(X)$ is the associated graded group built from the skeletal filtration of $K^{*}(X)$ built from the filtration of $X$. The differential

### v3

@@ 17. Clifford Algebras and Their Associated Lie Groups
Let’s look at the first few of these. For ${\mathcal{C}}_{0}$ we are handed $\mathbb{R}$} itself. The algebra ${\mathcal{C}}_{1}$ is formally given by
${\mathcal{C}}_{1}\cong({\mathbb{R}}\cdot 1)\oplus({\mathbb{R}}\cdot e_{1})$
which is easily seen to be isomorphic to the complex numbers $\mathbb{C}$}. Similarly, ${\mathcal{C}}_{2}$ is formally given by
${\mathcal{C}}_{2}\cong({\mathbb{R}}\cdot 1)\oplus({\mathbb{R}}\cdot e_{1})\oplus({\mathbb{R}}\cdot e_{2})\oplus({\mathbb{R}}\cdot e_{1}e_{2})$
which is isomorphic to the quaternions $\mathbb{H}$}.
@@
    ${\mathcal{C}}_{k+8}\,\cong\,{\mathcal{C}}_{k}\otimes_{{\mathbb{R}}}M_{16}({\mathbb{R}})\qquad\qquad{\mathcal{C}}^{c}_{k+2}\,\cong\,{\mathcal{C}}^{c}_{k}\otimes_{{\mathbb{C}}}M_{2}({\mathbb{C}})$
@@
Each Clifford algebra $A$ has a Lie group $Lie(A)\subset A$ associated to it, namely the elements $a\in A$ that are “unitary” in the sense that $aa^{*}=a^{*}a=1$. We write
@@
We may realize the Lie groups ${\mathcal{L}}_{j}$ of the Clifford algebras in a different way. Fix some large index $t$ and integer $k$. Fix a faithful $C^{*}$-representation
@@
$\big<O_{16t},\,\,\,J_{1},\dots J_{k}\big$>
denote the subgroup of $O_{16t}$ consisting of all of the matrices that commute with $J_{1},\dots J_{k}$. Then we have the following result.
@@
$U_{8t}$ $\cong\big<O_{16t},\,\,\,J_{1}\big$>
$Sp_{4t}$ $\cong\big<O_{16t},\,\,\,J_{1},J_{2}\big$>
$Sp_{2t}\times Sp_{2t}$ $\cong\big<O_{16t},\,\,\,J_{1},J_{2},J_{3}\big$>
$Sp_{2t}$ $\cong\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4}\big$>
$U_{2t}$ $\cong\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5}\big$>
$O_{2t}$ $\cong\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6}\big$>
$O_{t}\times O_{r}$ $\cong\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6},J_{7}\big$>
$O_{t}$ $\cong\big<O_{16t},\,\,\,J_{1},J_{2},J_{3},J_{4},J_{5},J_{6},J_{7},J_{8}\big$>
The sequence may be extended in each direction when $t$ is a sufficiently large power of $2$. The analogous result holds in the complex setting as well.
@@
$O_{1}\hookrightarrow U_{1}\hookrightarrow Sp_{1}\hookrightarrow Sp_{1}\times Sp_{1}\hookrightarrow Sp_{2}\hookrightarrow U_{4}\hookrightarrow O_{8}\hookrightarrow O_{8}\times O_{8}\hookrightarrow O_{16}$ (17.1)
@@
$O_{16}\hookleftarrow U_{8}\hookleftarrow Sp_{4}\hookleftarrow Sp_{2}\times Sp_{2}\hookleftarrow Sp_{2}\hookleftarrow U_{2}\hookleftarrow O_{2}\hookleftarrow O_{1}\times O_{1}\hookleftarrow O_{1}$ (17.2)
@@
from Sequence (17.2). So the groups appear in opposite order. What is really neat is that the maps are the same - just re-ordered! We have inserted the numbering of the maps simply to facilitate comparison. The same thing happens in the complex setting, but there are only two groups involved, and so it passes without notice. We want to refer to these two different points of view.
@@ 18. Clifford Algebras and Bott Periodicity: the Real Case
    There is a weak homotopy equivalence [^76] A based map $f:X\to Y$ is a weak homotopy equivalence if the induced map $f_{*}:\pi_{*}(X)\to\pi_{*}(Y)$ is an isomorphism. Every homotopy equivalence is a weak homotopy equivalence but not conversely.
@@
    Taking $A={\mathbb{R}}$ and $r=7$ in Part (3) above, there is a weak homotopy equivalence
@@
Please note that this strongly ties Clifford algebras, or, more precisely, their Lie groups, to the Homotopy Road Map, as shown in Table 11.2. Rather than regarding the first column as coming to us from Bott Periodicity, we may think of the column as simply presenting yet a different look at the Clifford and Lie algebra perspectives.
@@ 19. Clifford Algebras and Bott Periodicity: the Complex Case
Suppose that $A$ is a unital complex Banach algebra. Here’s some notation that we will use in this section:
@@
    Taking $A={\mathbb{C}}$, there is a weak homotopy equivalence
@@ 20. The Basics: Symmetric Spaces and a Timely Example
Suppose that $M$ is a connected Riemann manifold and let $p\in M$. There is an exponential map
$\exp_{p}:T_{p}M\to M.$
We may choose a small open ball $B_{p}$ about the origin of $T_{p}M$ and label its image as
$V_{p}\equiv\exp_{p}(B_{p})\subset M.$
If $B_{p}$ is small enough then the map
$\exp_{p}:B_{p}\longrightarrow V_{p}$
@@
$V_{p}\overset{\exp_{p}^{-1}}{\longrightarrow}B_{p}\overset{u\to-u}{\longrightarrow}B_{p}\overset{\exp_{p}}{\longrightarrow}V_{p}.$
@@
A Lie algebra $\mathfrak{g}$ over a field $\mathbb{F}$} is an $\mathbb{F}$}-vector space $\mathfrak{g}$ together with a map
@@
For example, take $M_{k}({\mathbb{F}})$ with [A,B]=ABBA[A,B]=AB-BA. [^80] Everyone should do the following exercise once in their lives. Suppose that $\mathfrak{g}=M_{k}({\mathbb{F}})$ with [A,B]=ABBA[A,B]=AB-BA. Then check explicitly that $\mathfrak{g}$ satisfies the Jacobi identity by calculating $[A,[B,C]]+[B,[C,A]]+[C,[A,B]]$ for $A,B,C\in M_{k}({\mathbb{F}})$.
@@
Now, what is the relationship between Lie groups and Lie algebras? Here are the facts:
@@
    If $G$ is a Lie group then $Lie(G)$ is a Lie algebra, with $dim(G)$ (as a compact smooth manifold) equal to $dim(Lie(G))$ (as a real vector space.)
@@
    If we restrict attention to connected, simply connected828282i.e. with $\pi_{1}(G)=0$. Lie groups, then there is a bijective correspondence
@@ A Timely Example
  5. (5)
    If instead of using $\Theta$ we used $Q$, defining
    $T^{\prime}(A)=QAQ^{\dagger}$
    then
    $T^{\prime}:\mathfrak{su}_{4}\longrightarrow\mathfrak{su}_{4}$
    and $T^{2}=-1$.
@@ 20.1. Topology and Physics, by Nadav Orion
for the joint system (both spaces with the usual inner product). A typical Hamiltonian for such a system contains both terms that affect a single qubit ($\sigma_{j}\otimes I$ and $I\otimes\sigma_{k}$) and terms that describe interactions between the two ($\sigma_{j}\otimes\sigma_{k}$), such that the set of Hamiltonians is all of the Hermitian $4\times 4$ matrices, which is the Lie algebra $\mathfrak{u}_{4}$. The evolution operator space is thus $U_{4}$, a well-known compact complex symmetric space.
@@ 21. Quantum Symmetry
Suppose that $G$ is a compact Lie group and $K$ is a Lie subgroup of $G$. Then the Lie algebra of $G$ decomposes as real vector spaces
@@
There are, however, other possibilities, the first of which being time-reversal symmetry $T$. Possessing this symmetry means that the system would look the same, even at the atomic level, if time were reversed. So in general, as we saw in the previous section, we can write time-reversal symmetry as $T$ with $T^{2}=I$ or $T^{2}=-I$ and the custom is to write 0 if it is not present. The Timely Example was indeed an example of time-reversal symmetry with $T^{2}=I$.
@@
It is possible to have a system that has neither time reversal nor charge conjugation symmetry, but it’s symmetrical under the combination of time reversal and charge conjugation. We also may have $S=CT$ with $S^{2}=1$ .
@@ C BdG 0 -1 0 $\mathbf{Sp}\,\,\,\cong\,\,\,\Omega^{4}\mathbf{O}$ ${\mathcal{R}}_{5}$ 0
we see at once that $KO$-theory and $K$-theory are intimately related to this story. The homotopy properties of the ${\mathcal{R}}_{n}$ are detected by $KO$ and hence connected to real vector bundles. This allows topological properties of the various symmetries to be detected and utilized. Much current physics research in topological insulators and related matters uses $KO^{*}$, and this was one of the main motivations for writing this monograph.
@@
22. Connecting Real and Complex $K$-theory
@@
    The product structure on the ring $KO_{*}({\mathbb{R}})$;
@@
    The complexification map $c\colon KO_{*}(A)\rightarrow K_{*}(A_{\mathbb{C}})$ is defined in Section 6.
@@
    But this projection is unitarily equivalent to $\left(\begin{smallmatrix}a&-b\\ b&a\end{smallmatrix}\right)$, by conjugating by $u=\left(\begin{smallmatrix}1&0\\ 0&-1\end{smallmatrix}\right)$. It follows that $r\psi[p]=r[p]$.
@@
That is, $KO_{*}({\mathbb{R}})$ is the quotient of the graded polynomial ring over the integers generated by $\eta,\xi,\beta_{O},$ and $\beta_{0}^{-1}$ modulo the ideal generated by the relations specified above.
@@
$\cdots\rightarrow KO_{n}(A)\xrightarrow{~\eta~}KO_{n+1}(A)\xrightarrow{~c~}K_{n+1}(A_{\mathbb{C}})\xrightarrow{~r\beta^{-1}}KO_{n-1}(A)\rightarrow\cdots$ (22.1)
@@
The following corollary provides a very important and, perhaps, unexpected connection between $KO$ and $K$.
@@
Consider the map $f_{*}$ applied to the short exact sequence of Equation (22.1).
@@
If $KO_{n}(A)=0$ for all $n$, then the long exact sequence (22.1) immediately implies that $K_{n}(A_{\mathbb{C}})=0$ for all $n$. Conversely, suppose that $K_{n}(A_{\mathbb{C}})=0$ for all $n$, then the long exact sequence (22.1) implies that $\eta_{n}$ is an isomorphism for all $n$. However, the relation $\eta^{3}=0$ then implies that $KO_{n}(A)=0$ for all $n$. ∎
@@
    If $A$ is a real filtered $C^{*}$-algebra, then there is an induced filtration on the complexification $A_{\mathbb{C}}$ given by
    $(A_{0})_{\mathbb{C}}\subset(A_{1})_{\mathbb{C}}\subset(A_{2})_{\mathbb{C}}\subset\dots$
    The complexification map $A\overset{c}{\longrightarrow}A_{\mathbb{C}}$ respects the filtrations on $A$ and $A_{\mathbb{C}}$ so that the map $c$ induces a natural transformation of spectral sequences
    $E_{p,q}^{r}(A)\overset{c}{\longrightarrow}E_{p,q}^{r}(A_{\mathbb{C}})$
    which is given by
    $KO_{i}(A_{p}/A_{p-1})\overset{c}{\longrightarrow}K_{i}((A_{p})_{\mathbb{C}}/(A_{p-1})_{\mathbb{C}})$
    on the $E^{1}$ pages and which converges to
    $KO_{i}(A)\overset{c}{\longrightarrow}K_{i}(A_{\mathbb{C}})\;.$
@@
Suppose that $X$ is a finite simplicial complex868686imagine a CW-complex with finitely many cells attached with skeleta [^87] For CW complexes, let $X_{0}$ be the finite set of points you start with, $X_{1}$ is what you have after adding all of the one-dimensional cells, etc. $X_{j}$. Let $\{x_{o}\}$ denote the space with a single point. Then there is a spectral sequence $E_{r}^{p,q}$ ($r\geq 1$, $\infty<p,q<+\infty$), with
$E_{1}^{p,q}\cong C^{p}(X;KO^{q}(\{x_{o}\})$
($p$- cochains taking values in $KO^{q}(\{x_{o}\}$ ), with $d^{1}$ usual coboundary operator, and
$E_{2}^{p,q}\cong H^{p}(X;KO^{q}(\{x_{o}\})$
$E_{\infty}^{p,q}\cong G_{p}KO^{p+q}(X).$
where $G_{p}KO^{p+q}(X)$ is the associated graded group built from the skeletal filtration of $KO^{*}(X)$ built from the filtration of $X$. The differential